OpenStudy (anonymous):

ln((x)/(x-1))+ln((x+1)/(x))-ln(x^2-1)

OpenStudy (anonymous):

\[change \ln (x^2-1) into \rightarrow \ln (x+1)(x-1)\rightarrow than \ln (x+1)+\ln (x-1)\]

OpenStudy (anonymous):

\[\ln (x+1)/x \rightarrow \ln (x+1)-\ln x\]

OpenStudy (anonymous):



OpenStudy (anonymous):

add them all and u will get final answer